Diseases tend to spread according to the exponential growth model. In the early days of AIDS, the growth factor (i.e. common ratio; growth multiplier) was around 1.9. In 1983, about 1600 people in the U.S. died of AIDS. If the trend had continued unchecked, how many people would have died from AIDS in 2006?
people

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

growth factor is 1.9.
that means that the number of people who got aids is 1.9 times as many as the year before.
in 1983, 1600 people in the united states died of aids.
with that annual growth factor, the number of people who would have died of aids in the us in 2006 would have been 1600 * 1.9 ^ (2006 - 1983) = 1600 * 1.9 ^ 23 = 4,125,274,047.
that's roughly 4 billion.
it's not very realistic, since the total population was only about 300,000,000 somewhere around 2006.


Answer by ikleyn(52864)

.

Correction to the post by @Theo.

"During 2005 alone, an estimated 2.8 million persons died from AIDS" - in the entire world.



Cited from this source
https://www.cdc.gov/mmwr/preview/mmwrhtml/mm5531a1.htm